OpenStudy (anonymous):

What would be the most logical first step to solve this quadratic equation? (x+5)^2=36

OpenStudy (anonymous):

Root both sides. Noting if y^2 = 36 , then y = ± 6

OpenStudy (anonymous):

Of course, if you'd asked ' which would be the most fun first step? ' , then that would be the last on my list.

OpenStudy (anonymous):

The most logical (and fast) would be rooting both sides: sqrt((x+5)²)=sqrt(36) x+5 = 6 x = 6-5 x = 1

OpenStudy (anonymous):

Martyr, you're wrong.

OpenStudy (anonymous):

Oh, God, my bad. x+5 = ±6 then x would be ±1

OpenStudy (anonymous):

Close.

OpenStudy (anonymous):

x actually equals 1 or -11.

OpenStudy (anonymous):

I can't believe I did it wrong again. So sorry about that. Thanks INewton.